Suppose Hilda’s Handbags makes large handbags and small handbags. Hilda sold 70 large bags for $450 each and 25 small bags for $150 each. If Hilda had total costs of $20,000, then Hilda’s profit would be _____ dollars.

*Enter your answer as an integer amount only, using no units or commas.
Answer #1

Ans) 15250

Total revenue = 70* 450 + 25* 150

= 31,500 + 3750 = 35,250

Profit = 35,250 - 20,000 = 15250